OpenStudy (anonymous):

What is the solution to this system of linear equations? 3x + y = -15 2x – 3y = 23

sam (.sam.):

3x + y = -15........multiply -3 2x – 3y = 23 -9x-3y=45 2x – 3y = 23 Elimination, -11x =22 x=-2 y=-9

OpenStudy (anonymous):